speed of boat in still water is 18km/h it takes 1/2hour extra in going 12km upstream instead of going the same distance downstream find speed of stream
Answer by stanbon(75887)

speed of boat in still water is 18km/h it takes 1/2hour extra in going 12km upstream instead of going the same distance downstream find speed of stream
------------------
Upstream DATA:
dist = 12 km ; time = t + (1/2) hrs ; rate = d/t = 12/(2t+1)/2 = 24/(2t+1) kph
-----------------------
Downstream DATA:
dist = 12 km ; time = t hrs ; rate = 12/t kph
--------
Equations:
Upstream:: 24/(2t+1) = 18-s
Downstream:: 12/t = 18+s
--------------------------------
Modify and solve:
s = [12/t]-18
Substitute::
24/(2t+1) = 18-[(12/t)-18]
-----
24/(2t+1) = 36-(12/t)
24 = 36(2t+1)-(12/t)(2t+1)
----
24 = 72t + 36 - [24+12/t)]
-12 = 72t - 24 - 12/t
12/t - 72t +12 = 0
----
1/t - 6t + 1 = 0
-6t^2 +t + 1 = 0
(-2t+1)(3t+1) =
Positive solution:
t = 1/2
------
Solve for "s":
s = [12/t]-18
stream speed = (12/(1/2) - 18 = 24-18 = 6 km/hr
------------------
Cheers,
Stan H.
-----------------